### v1

# Flows of vector fields and the Kalman Theorem

Abstract.
We investigate the relations between the Kalman Theorem and the Chow-Rashevskiĭ Theorem or, more precisely, the general theory of flows tangent to non-integrable distributions. The main results consist of two proofs of the Kalman Theorem, which are alternative to the most common ones and which enlighten the above relations and stimulate generalisations in various directions.
Key words and phrases:
Controllability; Flow of vector fields; Kalman Theorem; Chow-Rashevskiĭ Theorem; Linear systems
2020 Mathematics Subject Classification:
34C99; 95B05; 93C05
1. Introduction
Consider a controlled dynamical system, whose states are represented by the points $q=(q^{i})_{1\leq i\leq n}$ in the affine space $\mathscr{Q}:=\mathbb{R}^{n}$ and with evolutions given by the solutions of a system of linear first order equations of the form
$\dot{q}=Aq+Bu$ (1.1)
for some non zero constant matrices $A=(A^{i}_{j})$ and $B=(B^{i}_{a})$ and in which the controls $u(t)$ are measurable functions taking values in a bounded set $\mathscr{K}\subset\mathbb{R}^{m}$. For a fixed $\overline{q}\in\mathscr{Q}$, denote by $\mathcal{C}_{\overline{q}}$ the set of all points $q\in\mathscr{Q}$ that are controllable to $\overline{q}$, i.e. for which there is a control $u(t)$ which determines a solution $q(t)$ to (1.1) that starts from $q$ and reaches $\overline{q}$ in a finite time. The famous Kalman Theorem is an important criterion to determine whether the controllable points to the origin $0_{\mathbb{R}^{n}}$ contains an open set. It states that
Theorem (Kalman). If $\mathscr{K}$ contains a neighbourhood of $0_{\mathbb{R}^{m}}$, the set $\mathcal{C}_{\overline{q}=0_{\mathbb{R}^{n}}}$ contains a neighbourhood of $0_{\mathbb{R}^{n}}$ if and only if the space $V\subset\mathbb{R}^{n}$, which is spanned by the vectors
$B_{1},\ldots,B_{m},\ \ A{\cdot}B_{1},\ldots,A{\cdot}B_{m},\ \ (A{\cdot}A){\cdot}B_{1},(A{\cdot}A){\cdot}B_{1}\ldots,\ (A{\cdot}A){\cdot}B_{m},\ldots\\ \ldots,\ \ \left(A{\cdot}\ldots{\cdot}A\right){\cdot}B_{1},\left(A{\cdot}\ldots{\cdot}A\right){\cdot}B_{2},\ldots,\left(A{\cdot}\ldots{\cdot}A\right){\cdot}B_{m},\ \ \ldots\ ,$ (1.2)
has the maximal dimension $n$ (here, we denote by $B_{a}$ the columns of $B$).
The proof can be found in many textbooks (as for instance [1, 3]), and it essentially uses just results in Linear Algebra and Convex Geometry.
This maximal dimension condition is very similar to the condition on the Lie bracket generated spaces, which appear in the corollary of the Chow-Rashevskiĭ Theorem concerning the local controllability of systems of the form $\dot{q}=\sum_{a=0}^{m}X_{a}(q)u^{a}$. In fact, the relation between those theorems is much more than a mere likeness, since
  • (a)
    The vector space $V$ generated by the vectors (1.2) coincides with the space that is generated by the Lie brackets at $q=0$ of an appropriate collection of $m+1$ vector fields $X_{a}:\mathscr{Q}\to\mathbb{R}^{n}$ (and this is exactly the space considered in the statement of the Chow-Rashevskiĭ Theorem (see §3.2 below)), namely of a collection of vector fields, among which the components of first vector field $X_{0}$ are linear functions of the coordinates of $\mathscr{Q}$, while the components of all the others are constant functions;
  • (b)
    The sufficiency parts of both results, i.e. of the Kalman Theorem and of the mentioned corollary of the Chow-Rashevskiĭ  Theorem, can be derived from a very general criterion by Sussmann ([13]; see also [3, Thm. 3.29 & Ex.3.3]);
  • (c)
    A couple of results by Hermann [6] and Nagano [10] provide a common argument for proving the necessity parts of both theorems (also on this regard, see e.g. [1, 3]).
Motivated by these facts, there is a common belief that the Kalman Theorem is essentially equivalent to the above quoted corollary of the Chow-Rashevskiĭ Theorem. Since it is not hard to check that “the Kalman Theorem yields such a corollary of the Chow-Rashevskiĭ Theorem” (see e.g. §4 below), the existence or not of an equivalence between those two theorems depends on whether or not the Chow-Rashevskiĭ Theorem implies the Kalman Theorem in a direct way. The aim of this paper is precisely to clarify to which extent such an implication exists and hence whether the above expected equivalence is actually correct.
Our achievements consist of two (at best of our knowledge, new) proofs for the Kalman Theorem that answer the above question and that we believe can be used also to determine convenient generalisations of the Kalman Theorem. The first proof is short, direct and makes fully explicit a detailed relation between the (above described corollary of the) Chow-Rashevskiĭ  Theorem and the Kalman Theorem. It shows that the first implies the second on the basis of the following simple (but crucial) fact: if $\mathscr{K}$ contains a neighbourhood of the origin – and thus, in particular, a convex symmetric neighbourhood of the $0_{\mathbb{R}^{n}}$ – then also the controllable set $\mathcal{C}_{q=0_{\mathbb{R}^{n}}}$ of (1.1) contains a convex set (see (1) of Proposition 5.1, below). We guess that this new proof admits generalisation for other kinds of (possibly non-linear) control systems, provided that similar convexity properties for the controllable sets can be established. On this regard, we would like to mention that investigations, involving ideas very close to those of our first proof, can be found in [8], where interesting extensions of Kalman criterion are reached.
Our second proof is much longer than the first, it is based on a discussion on the higher dimensional space =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$ and it does not involve in a direct way the Chow-Rashevskiĭ  Theorem – it just uses a few results on flows, which are properties that can be taken as ancestors of the Chow-Rashevskiĭ  Theorem. However, these drawbacks are balanced by the fact that our second proof makes no use of any information on convexity properties of the controllable sets. This makes this second proof even more appropriate than the first for generalisations to non-linear control systems. In fact, generalisations of this kind have been established in [5], where new local controllability criteria are given and diverse controllability problems, for which all previously known criteria are inconclusive, are solved.
The paper is structured as follows. After section §2, in which we give a few preliminaries and introduce some convenient notation, in §3 we recall the statements of the Chow-Rashevskiĭ and the Kalman Theorem, on which our discussion is based. In §4 we prove that, for the collections of $m+1$ vector fields $X_{\alpha}$ of the above described form, the property determined by the Chow-Rashevskiĭ Theorem can be inferred directly from the Kalman Theorem. In §5 and §6 there are our main results, i.e. the mentioned couple of new proofs for the Kalman Theorem, one based on the Chow-Rashevskiĭ Theorem and the other on general properties of flows.
Throughout this paper we adopt the Einstein convention on summations.
2. Preliminaries
2.1. Vector fields, Lie brackets and flows
We recall that a $\mathcal{C}^{k}$ vector field on $\mathscr{U}\subset\mathbb{R}^{n}$ is an application in $\mathcal{C}^{k}(\mathscr{U},\mathbb{R}^{n})$, which maps each $x=(x^{i})_{1\leq i\leq n}\in\mathscr{U}$ into a vector $X(x)=(X^{1}(x),\ldots,X^{n}(x))$ applied to the point $x$. We also recall that each $\mathcal{C}^{k}$ vector field $X=(X^{i})_{1\leq i\leq n}$ can be identified with the first order differential operator on real functions $X(f):=X^{i}\frac{\partial f}{\partial x^{i}}$. Under this identification, the Lie bracket $[X,Y]$ between two $\mathcal{C}^{k}$ vector fields $X,Y$ on $\mathscr{U}$, $k\geq 1$, can be defined as the $\mathcal{C}^{k-1}$ vector field on $\mathscr{U}$ identified with the first order differential operator
$[X,Y](f):=X(Y(f))-Y(X(f))=\left(X^{j}\frac{\partial Y^{i}}{\partial x^{j}}-Y^{j}\frac{\partial X^{i}}{\partial x^{j}}\right)\frac{\partial f}{\partial x^{i}}=\\ =\left(X(Y^{i})-Y(X^{i})\right)\frac{\partial f}{\partial x^{i}}\ .$
Given a $\mathcal{C}^{k}$ vector field $X$ on $\mathscr{U}\subset\mathbb{R}^{n}$, a corresponding (local) flow on a relatively compact open subset $\widetilde{\mathscr{U}}\subset\mathscr{U}$ is a map $\Phi^{X}:(-\varepsilon,\varepsilon)\times\widetilde{\mathscr{U}}\longrightarrow\mathbb{R}^{n}$, which is constructed as follows for a sufficiently small $\varepsilon>0$ For any $x\in\widetilde{\mathscr{U}}$, let us denote by $\gamma^{(x)}(s)$, $s\in(-\varepsilon^{(x)}_{1},\varepsilon^{(x)}_{2})$, the unique maximal solution to the differential problem $\dot{\gamma}^{(x)}(s)=X(\gamma^{(x)}(s))$, $\gamma^{(x)}(0)=x$. Given $0<\varepsilon\leq\min_{q\in\widetilde{\mathscr{U}}}\{\varepsilon^{(x)}_{1},\varepsilon^{(x)}_{2}\}$, the flow $\Phi^{X}$ is the map defined by
$\Phi^{X}:(-\varepsilon,\varepsilon)\times\widetilde{\mathscr{U}}\longrightarrow\mathbb{R}^{n}\ ,\qquad\Phi^{X}(s,x):=\gamma^{(x)}(s)\ .$
By construction, for any fixed $s_{o}\in(-\varepsilon,\varepsilon)$ the map $\Phi^{X}_{s_{o}}:=\Phi^{X}(s_{o},\cdot):\widetilde{\mathscr{U}}\to\mathbb{R}^{n}$ is a diffeomorphism from $\widetilde{\mathscr{U}}$ onto $\Phi^{X}_{s_{o}}(\widetilde{\mathscr{U}})$, with inverse $\Phi^{X}_{-s_{o}}=\Phi^{-X}_{s_{o}}$, and with
$\frac{\partial\Phi^{X}(s,x)}{\partial s}\bigg{|}_{s=0}=X(x)\qquad\text{for any}\ x\in\widetilde{\mathscr{U}}\ .$
Moreover, $\Phi^{X}_{s_{o}=0}=\operatorname{Id}_{\widetilde{\mathscr{U}}}$ and, whenever $s,s^{\prime}$ are such that $s+s^{\prime}\in(-\varepsilon,\varepsilon)$ and $\Phi^{X}_{s}(x),\Phi^{X}_{s^{\prime}}(x)\in\widetilde{\mathscr{U}}$,
$\Phi^{X}_{s+s^{\prime}}(x)=\Phi^{X}_{s}\big{(}\Phi^{X}_{s^{\prime}}(x)\big{)}=\Phi^{X}_{s^{\prime}}\big{(}\Phi^{X}_{s}(x)\big{)}\ .$
2.2. Pushed-forward vector fields and their flows
Let $X$ be a $\mathcal{C}^{k}$ vector field on an open set $\mathscr{U}\subset\mathbb{R}^{n}$ and $\varphi:\mathscr{U}\to\mathscr{V}=\varphi(\mathscr{U})\subset\mathbb{R}^{n}$ a diffeomorphism between $\mathscr{U}$ and $\mathscr{V}\subset\mathbb{R}^{n}$. The push-forward of $X$ by the map $\varphi$ is the vector field on $\mathscr{V}$
$\varphi_{*}(X)=\big{(}\varphi_{*}(X)^{1},\ldots\varphi_{*}(X)^{n}\big{)}$
corresponding to the linear differential operator defined by
$\varphi_{*}(X)(f)\big{|}_{q}:=X(f\circ\varphi)\big{|}_{\varphi^{-1}(q)}=X^{j}(\varphi^{-1}(q))\frac{\partial(f\circ\varphi)}{\partial q^{j}}\bigg{|}_{\varphi^{-1}(q)}=\\ =X^{j}(\varphi^{-1}(q))\frac{\partial\varphi^{i}}{\partial q^{j}}\bigg{|}_{\varphi^{-1}(q)}\frac{\partial f}{\partial q^{i}}\bigg{|}_{q}=\left(J(\varphi)^{i}_{j}X^{j}\right)\bigg{|}_{\varphi^{-1}(q)}\frac{\partial f}{\partial q^{i}}\bigg{|}_{q}\ ,$ (2.1)
where $J(\varphi)$ is the Jacobian matrix of $\varphi$. In other words, $\varphi_{*}(X)$ is the vector field on $\mathscr{V}$ with components $\varphi_{*}(X)^{i}(q)=\bigg{(}J(\varphi)^{i}_{j}X^{j}\bigg{)}(\varphi^{-1}(q))$.
The following fact is well known (see e.g. [14, Prop. 1.55]): For any two $\mathcal{C}^{k}$ vector fields $X$, $Y$, $k\geq 1$, on $\mathscr{U}$ and any $\mathcal{C}^{2}$ diffeomorphism $\varphi:\mathscr{U}\to\mathscr{V}\subset\mathbb{R}^{n}$ then $\varphi_{*}([X,Y])=\left[\varphi_{*}(X),\varphi_{*}(Y)\right]$. This and a direct calculation imply that the following relation between flows holds at all points where both sides are defined:
$\Phi^{X}_{s}\circ\Phi^{Y}_{t}=\Phi_{t}^{(\Phi^{X}_{s})_{*}(Y)}\circ\Phi^{X}_{s}\ .$ (2.2)
2.3. Piecewise regular curves and their compositions
Given a manifold $\mathscr{N}$, by parameterised regular curve we mean a $\mathcal{C}^{1}$ map $\gamma:[a,b]\subset\mathbb{R}\longrightarrow\mathscr{N}$ with nowhere vanishing velocity $\dot{\gamma}(t)$ and which is an homeomorphism between $[a,b]$ and the image $\gamma([a,b])$ of the map. The image $\gamma([a,b])\subset\mathscr{N}$ is called (non-parameterised) regular curve. Two parameterised regular curves $\gamma(t)$, $\widetilde{\gamma}(s)$ with the same image are called consonant if one is determined from the other via a change of parameter $t=t(s)$ with $\frac{dt}{ds}>0$ at all points. Consonance is clearly an equivalence relation and an orientation of a regular curve is a choice of one of the only two possible equivalence classes of its parameterisations. An oriented curve is a regular curve with an orientation. In the following, we denote it by means of one of the consonant parameterisations $\gamma(t)$ of the chosen equivalence class.
Given two oriented curves $\gamma_{1}(t)$, $\gamma_{2}(s)$, $t\in[a,b]$, $s\in[a^{\prime},b^{\prime}]$, with $\gamma_{1}(b)=\gamma_{2}(a^{\prime})$, their union is called (oriented) composition and we denote it by $\gamma_{1}\ast\gamma_{2}$. The oriented regular curves $\gamma_{1}$, $\gamma_{2}$ that give the composition $\gamma_{1}\ast\gamma_{2}$ are called regular arcs of the composition. In a similar way we define the (oriented) composition of a finite number of oriented regular curves, each of them sharing its final endpoint with the initial endpoint of the succeeding one. The subsets of $\mathscr{N}$, which are compositions of a finite number of oriented regular curves, are called piecewise regular (oriented) curves.
3. The Chow-Rashevskiĭ Theorem and the Kalman Theorem
3.1. Orbits of sets of vector fields and the Chow-Rashevskiĭ Theorem
Let =$\mathscr{F}=\{X_{A}\}_{A\in\mathscr{J}}$ be a family of $\mathcal{C}^{\infty}$ vector fields on an open subset $\mathscr{U}\subset\mathbb{R}^{n}$, indexed by the elements of a (possibly uncountable) set $\mathscr{J}$, and denote by $\widehat{\mathscr{F}}$ the space of all finite linear combinations with constant coefficients of the vectors in $\mathscr{F}$, i.e.
^=:=$\widehat{\mathscr{F}}=\langle\mathscr{F}\rangle:=\bigg{\{}Y=\lambda_{A_{1}}X_{A_{1}}+\ldots+\lambda_{A_{N}}X_{A_{N}}\ ,\ \lambda_{A_{i}}\in\mathbb{R}\ ,\ N\in\mathbb{N}\bigg{\}}\ .$
The orbit by $\mathscr{F}$ of a point $x_{o}\in\mathbb{R}^{n}$ is the subset of $\mathbb{R}^{n}$
$\operatorname{Orb}^{\mathscr{F}}(x_{o})=\bigg{\{}\ x\in\mathscr{U}\ :\ x=\Phi^{Y_{1}}_{s_{1}}\circ\ldots\circ\Phi^{Y_{N}}_{s_{N}}(x_{o})\ ,\ \ N\in\mathbb{N}\ ,\ s_{i}\in\mathbb{R}\ ,\ Y_{i}\in\widehat{\mathscr{F}}\ \bigg{\}}\ .$ (3.1)
The Lie span of $\mathscr{F}$ at the point $x_{o}$ is the linear subspace of $T_{x_{o}}\mathbb{R}^{n}=\mathbb{R}^{n}$ defined by
$\operatorname{Lie}(\mathscr{F})_{x_{o}}:=\bigg{\{}\ v=[Y_{1},[Y_{2},\ldots[Y_{N-2},[Y_{N-1},Y_{N}]]\ldots]]\big{|}_{x_{o}}\ ,\ N\in\mathbb{N}\ ,\ Y_{j}\in\widehat{\mathscr{F}}\bigg{\}}\ .$
This notation being fixed, we can state Sussmann’s refined and improved version of the Chow-Rashevskiĭ Theorem as follows (see also [4, 5]).
Theorem 3.1 (Chow-Rashevskiĭ-Sussmann [2, 11, 12]).
For any family $\mathscr{F}$ of $\mathcal{C}^{\infty}$ vector fields on an open subset $\mathscr{U}\subset\mathbb{R}^{n}$ and for any $x_{o}\in\mathscr{U}$, the corresponding set $\operatorname{Orb}^{\mathscr{F}}(x_{o})$ is a connected immersed submanifold of $\mathbb{R}^{n}$ of dimension greater than or equal to $\dim\operatorname{Lie}(\mathscr{F})_{x_{o}}$. Moreover, if the vector fields are real analytic, then $\operatorname{Orb}^{\mathscr{F}}(x_{o})$ contains a neighbourhood of $x_{o}$ if and only if $\dim\operatorname{Lie}(\mathscr{F})_{x_{o}}=n$.
3.2. The Kalman Theorem
Consider the system (1.1), which, for brevity, in the following will be called the linear system $(A,B)$. Given a measurable control $u_{o}:[0,T]\to\mathscr{K}$, we denote by $q^{(A,B|q_{o},u_{o})}:[0,T]\to\mathscr{Q}$ the unique absolutely continuous map which solves the equations with control $u_{o}$ and initial condition $q_{o}$, that is the solution to
$\begin{array}[]{ll}\dot{q}(t)=Aq(t)+Bu_{o}(t)\ \text{for a.e.}\ t\in[0,T]\ ,\\ q(0)=q_{o}\ .\end{array}$ (3.2)
For any choice of points $\overline{q},q_{o}\in\mathscr{Q}$ we may consider the following associated subsets of $\mathscr{Q}$:
  • The set of the states that are controllable to $\overline{q}$, i.e. the set of $q\in\mathscr{Q}$ from which the controlled system can start to reach the prescribed final state $\overline{q}$ in a finite time $T\geq 0$;
  • The set of all states that are reachable from $q_{o}$, i.e. the set of $q\in\mathscr{Q}$ that can be reached in finite time $T\geq 0$ starting from the prescribed initial state $q_{o}$.
We denote such two sets, respectively, by
$\mathcal{C}^{(A,B)}_{\overline{q}}:=\bigg{\{}q_{o}\in\mathbb{R}^{n}\ :\ q^{(A,B|q_{o},u_{o})}(T)=\overline{q}\ ,\ \text{for some}\ T\geq 0\ \text{and}\ u_{o}:[0,T]\to\mathscr{K}\ \text{meas.}\bigg{\}}\ ,$. } , (3.3)
$\mathscr{O}^{(A,B)}_{q_{o}}:=\bigg{\{}\overline{q}\in\mathbb{R}^{n}\ :\ q^{(A,B|q_{o},u_{o})}(T)=\overline{q}\ ,\ \text{for some}\ T\geq 0\ \text{and}\ u_{o}:[0,T]\to\mathscr{K}\ \text{meas.}\bigg{\}}\ .$. } . (3.4)
The subsets of these two sets, given by the points that are controllable to $\bar{q}$ or reachable from $q_{o}$, respectively, by means of piecewise constant controls are denoted by $\mathcal{C}^{(A,B)\text{p.c.}}_{\overline{q}}$ and $\mathscr{O}^{(A,B)\text{p.c.}}_{q_{o}}$, respectively.
Due to the linearity of the system (1.1) and the fact that the matrices $A$ and $B$ are constant, the set of controllable states and the set of reachable states are related each other as follows. If we set $\mathbb{A}:=-A$, $\mathbb{B}=-B$, one can directly see that for any $q\in\mathscr{Q}=\mathbb{R}^{n}$,
$\mathscr{O}^{(A,B)}_{q}=\mathcal{C}^{(\mathbb{A},\mathbb{B})}_{q}\ ,\qquad\mathcal{C}^{(A,B)}_{q}=\mathscr{O}^{(\mathbb{A},\mathbb{B})}_{q}\ .$ (3.5)
Moreover, if the control set $\mathscr{K}$ is symmetric (i.e. such that $\mathscr{K}=-\mathscr{K}$) we also have that
$\mathcal{C}^{(A,B)}_{q}=\mathcal{C}^{(A,\mathbb{B})}_{q}=\mathscr{O}^{(\mathbb{A},B)}_{q}\ ,\qquad\mathscr{O}^{(A,B)}_{q}=\mathscr{O}^{(A,\mathbb{B})}_{q}=\mathcal{C}^{(\mathbb{A},B)}_{q}\ .$ (3.6)
Similar relations hold for the sets $\mathcal{C}^{(A,B)\text{p.c.}}_{q}$, $\mathscr{O}^{(\mathbb{A},B)\text{p.c.}}_{q}$, etc.
We are now ready to state the version of the Kalman Theorem, which we are going to use in all forthcoming discussion and which is well known to be fully equivalent to the statement in the Introduction. In a matter of clarity, in what follows the space $V$, generated by the vectors defined in (1.2), will be denoted by $V=\operatorname{Kalman}^{(A,B)}$.
Theorem 3.2 (Kalman Theorem).
If $\mathscr{K}\subset\mathbb{R}^{m}$ is a neighbourhood of $0_{\mathbb{R}^{m}}\in\mathbb{R}^{m}$, the set $\mathcal{C}^{(A,B)\text{\rm p.c.}}_{\overline{q}=0}$ contains a neighbourhood of $0_{\mathbb{R}^{n}}\in\mathbb{R}^{n}$ if and only if $\dim\operatorname{Kalman}^{(A,B)}=n$.
The equality $\dim\operatorname{Kalman}^{(A,B)}=n$ is commonly called Kalman condition.
4. The Kalman Theorem implies the corollary of the Chow-Rashevskiĭ Theorem for sets of vector fields with either constant or linear components
Given a linear system $(A,B)$, let $\mathscr{F}^{(A,B)}$ be the finite set of real analytic vector fields $\mathscr{F}^{(A,B)}:=\{X_{0},X_{1},\ldots,X_{m}\}$ on $\mathbb{R}^{n}$, with $X_{\alpha}$ defined at each point $q$ by
$\begin{split}&\hskip 113.81102ptX_{0}(q):=(-A^{1}_{j}q^{j},\ldots,-A^{n}_{j}q^{j})=-A\,q\qquad\text{and}\\ &X_{1}(q):=(B_{1}^{1},\ldots,B_{1}^{n})\ ,\ X_{2}(q):=(B_{2}^{1},\ldots,B_{2}^{n})\ ,\ \ldots,\ X_{m}(q):=(B_{m}^{1},\ldots,B_{m}^{n})\ .\end{split}$ (4.1)
Note that, for $\alpha=1,\ldots,m$, the components of the vector field $X_{\alpha}$ are precisely the entries of the $\alpha-$th column of the matrix $B$. Finally, for any $\lambda=(\lambda^{i})\in\mathbb{R}^{m}$, let us set $Y^{(\lambda)}:=X_{0}+\lambda^{i}X_{i}$.
We want to show that using the Kalman Theorem in a direct and elementary way, one can prove that the orbit $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(0_{\mathbb{R}^{n}})$ contains a neighbourhood of the origin if and only if $\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{q=0}=n$.
Indeed, for any $\widehat{q}\in\mathbb{R}^{n}$ and $\lambda\in\mathbb{R}^{m}$, the integral curve $t\mapsto\Phi^{Y^{(\lambda)}}_{t}(\widehat{q})$ of the vector field $Y^{(\lambda)}$ coincides with the solution $q^{(-A,B|\widehat{q},u_{o})}(t)$ of the linear system $(-A,B)$ determined by the constant control $u_{o}(t)\equiv\lambda$. Hence if we consider the linear system $(A,B)$ with a control set $\mathscr{K}$ which is not only open, but also symmetric, by (3.5) and (3.6) for any fixed point $\overline{q}$, the corresponding set $\mathcal{C}^{(A,B)\text{p.c.}}_{\overline{q}}\bigg{(}=\mathscr{O}^{(-A,B)\text{p.c.}}_{\overline{q}}\bigg{)}$ coincides with the set of points of the form
$q(T)=\Phi^{Y^{(\lambda_{N})}}_{s_{N}}\circ\ldots\circ\Phi^{Y^{(\lambda_{1})}}_{s_{1}}(\overline{q})\ ,\qquad T=s_{1}+s_{2}+\ldots+s_{N}>0\ ,$
for some finite sequence of $m$-tuples $\lambda_{i}=(\lambda_{i}^{1},\ldots,\lambda^{m}_{i})$. Since each such a point is in $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(\overline{q})$, we conclude that $\mathcal{C}_{\overline{q}}^{(A,B)\text{\rm p.c.}}\subset\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(\overline{q})$. From this it follows immediately that if the linear system $(A,B)$ satisfies the Kalman condition, then, by the Kalman Theorem, not only $\mathcal{C}_{\overline{q}=0_{\mathbb{R}^{n}}}^{(A,B)\text{\rm p.c.}}$ contains a neighbourhood of $0_{\mathbb{R}^{n}}$, but also the orbit $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(0_{\mathbb{R}^{n}})$ contains a neighbourhood of the origin as desired.
We claim that the converse is true as well. The proof of this is indeed elementary and needs neither the Kalman Theorem nor other results nor convexity assumptions, except some standard linear algebra. For the sake of completeness, here are the details.
Assume that the linear system $(A,B)$ does not satisfy the Kalman condition, i.e. that the vector space $\operatorname{Kalman}^{(A,B)}$ has dimension $n^{\prime}<n$, and pick a set of $n^{\prime}$ linearly independent vectors among the generators (1.2), say $C_{1},\ldots,C_{n^{\prime}}$. Then consider a set of $n-n^{\prime}$ additional vectors, denoted by $C_{n^{\prime}+1}$, …, $C_{n}$ not belonging to $\operatorname{Kalman}^{(A,B)}$ , which complete the previous vectors to a basis for $\mathbb{R}^{n}$. Applying an appropriate linear change of coordinates, there is no loss of generality if we assume that $C_{1}=(1,0,\ldots,0)$, $C_{2}=(0,1,\ldots,0)$, …, $C_{n}=(0,0,\ldots,1)$ and that
$\operatorname{Kalman}^{(A,B)}=\langle C_{1},\ldots,C_{n^{\prime}}\rangle=\{q_{n^{\prime}+1}=\ldots=q_{n}=0\}.$
Note that, being each vector $C_{i}$, $1\leq i\leq n^{\prime}$, in $\operatorname{Kalman}^{(A,B)}$, the corresponding vector $A\cdot C_{i}$ is also in $\operatorname{Kalman}^{(A,B)}$. This implies that the matrix $A$ (in the considered new coordinates!) has the form
$A=\left(\begin{array}[]{cccccc}a_{1}^{1}&\ldots&a_{n^{\prime}}^{1}&a_{n^{\prime}+1}^{1}&\ldots&a_{n}^{1}\\ a_{1}^{2}&\ldots&a_{n^{\prime}}^{2}&a_{n^{\prime}+1}^{2}&\ldots&a_{n}^{2}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ a_{1}^{n^{\prime}}&\ldots&a_{n^{\prime}}^{n^{\prime}}&a_{n^{\prime}+1}^{n^{\prime}}&\ldots&a_{n}^{n^{\prime}}\\ 0&\ldots&0&a_{n^{\prime}+1}^{n^{\prime}+1}&\ldots&a_{n}^{n^{\prime}+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&\ldots&0&a_{n^{\prime}+1}^{n}&\ldots&a_{n}^{n}\end{array}\right)\ .$ (4.2)
In particular, for any $q\in\operatorname{Kalman}^{(A,B)}$, also the vector $A{\cdot}q$ is in $\operatorname{Kalman}^{(A,B)}$. It follows that the value at any point of $\operatorname{Kalman}^{(A,B)}=\{q_{n^{\prime}+1}=\ldots=q_{n}=0\}$ of a vector field in $\mathscr{F}^{(A,B)}$ is a vector which is tangent to $\operatorname{Kalman}^{(A,B)}$ and that any flow of a vector field in $\widehat{\mathscr{F}^{(A,B)}}$ maps the subspace $\operatorname{Kalman}^{(A,B)}$ into itself. We conclude that if the linear system $(A,B)$ does not satisfies the Kalman condition, then the orbit $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(0_{\mathbb{R}^{n}})$ is contained in $\operatorname{Kalman}^{(A,B)}\subsetneq\mathbb{R}^{n}$ and cannot contain any neighbourhood of the origin, as claimed.
In order to conclude the proof of the claim
$\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(0_{\mathbb{R}^{n}})$ contains a neighbourhood of $0_{\mathbb{R}^{n}}$ if and only if $\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{q=0}=n$
it remains to show that, for such a kind of set of vector fields, the Kalman condition is equivalent to the condition $\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{q=0}=n$. For this, it suffices to notice that, due to the constancy of the vector fields $X_{1},\ldots,X_{m}$, in the set of all iterated Lie brackets $[X_{i_{1}},[X_{i_{2}},\ldots,[X_{i_{N-1}},X_{i_{N}}]\ldots]]$, $N\geq 2$, the only non trivial brackets are those having the form
[…]
From this it follows immediately that $\operatorname{Kalman}^{(A,B)}=\operatorname{Lie}(\mathscr{F}^{(A,B)})\big{|}_{\overline{q}=0}$ and this immediately gives the desired equality between the two dimensions.
5. A proof of the Kalman Theorem, based on the Chow-Rashewskiĭ Theorem
We give now our first new proof of the Kalman Theorem, which shows how that theorem can be obtained as a corollary of the Chow-Rashevkiĭ Theorem.
First of all, we stress the fact that the necessity part of Theorem 3.2 (or, equivalently, the fact that if the Kalman condition does not hold, then the set $\mathcal{C}^{(A,B)}_{\overline{q}=0_{\mathbb{R}^{n}}}=\mathscr{O}^{(-A,B)}_{\overline{q}=0_{\mathbb{R}^{n}}}$ cannot contain any neighbourhood of the origin) is a consequence of the same linear algebra arguments considered at the end of §4. Indeed, if $\dim\operatorname{Kalman}^{(A,B)}=n^{\prime}<n$, by those arguments, there is no loss of generality if we assume that $\operatorname{Kalman}^{(A,B)}$ is the space $\operatorname{Kalman}^{(A,B)}=\{q^{n^{\prime}+1}=\ldots=q^{N}=0\}$, that $A$ has the form (4.2) and that all vectors $B_{\alpha}$, $1\leq\alpha\leq m$, are in the subspace $\operatorname{Kalman}^{(A,B)}$. If this is the case, one can immediately see that any solution of the linear system $(-A,B)$ that starts from $0_{\mathbb{R}^{n}}$ is constrained to stay in $\operatorname{Kalman}^{(A,B)}$, so that $\mathscr{O}^{(-A,B)}_{\overline{q}=0_{\mathbb{R}^{n}}}=\mathcal{C}^{(A,B)}_{\overline{q}=0_{\mathbb{R}^{n}}}\subset\operatorname{Kalman}^{(A,B)}$ does not contain any neighbourhood of the origin.
Second, we point out that, in order to prove the sufficiency part of the Kalman Theorem, there is no loss of generality if we assume that $\mathscr{K}$ is not only a neighbourhood of $0_{\mathbb{R}^{n}}$, but is also convex and symmetric. Indeed, if we use the notation $\mathcal{C}^{(A,B)\text{\rm p.c.}}_{\overline{q}=0_{\mathbb{R}^{n}}}$ and $\widetilde{\mathcal{C}^{(A,B)\text{\rm p.c.}}}_{\overline{q}=0_{\mathbb{R}^{n}}}$ for the sets of points that are controllable to $0_{\mathbb{R}^{m}}$ by means of controls in $\mathscr{K}$ and of controls in a subset $\widetilde{\mathscr{K}}$ of $\mathscr{K}$, respectively, it is clear that $\widetilde{\mathcal{C}^{(A,B)\text{\rm p.c.}}}_{\overline{q}=0_{\mathbb{R}^{n}}}\subset\mathcal{C}^{(A,B)\text{\rm p.c.}}_{\overline{q}=0_{\mathbb{R}^{n}}}$. If in addition $\mathscr{K}$ contains a neighbourhood of $0_{\mathbb{R}^{m}}$, then it certainly contains a subset $\widetilde{\mathscr{K}}$ which is not only a neighbourhood of $0_{\mathbb{R}^{m}}$, but it also convex and symmetric (a ball, for example). Hence if we prove that the Kalman condition is sufficient for $\widetilde{\mathcal{C}^{(A,B)\text{\rm p.c.}}}_{\overline{q}=0_{\mathbb{R}^{n}}}$ to contain a neighbourhood of the origin, we immediately get that the Kalman condition a fortiori is a sufficient condition for $\mathcal{C}^{(A,B)\text{\rm p.c.}}_{\overline{q}=0_{\mathbb{R}^{n}}}$ having that property.
Due this observation, the proof of the sufficiency part of the Kalman Theorem can be made under the stronger assumption that $\mathscr{K}$ is open, convex, symmetric and contains the origin. For such a proof we need the following
Proposition 5.1.
If $\mathscr{K}\subset\mathbb{R}^{m}$ is open, convex, symmetric and contains the origin, then:
  • (1)
    $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$ is symmetric and convex;
  • (2)
    If $L\subset\mathbb{R}^{n}$ is the smallest (with respect to inclusion) linear subspace containing $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$, then $\dim L\geq 1$ and $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$ contains a neighbourhood of 00 of the topology of $L$ (considered as a topological subspace of $\mathbb{R}^{n}$);
  • (3)
    $\dim L=\dim\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(0)\geq\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{\overline{q}=0}$.
Proof.
(1) is a well known property of linear system, which can be easily checked by just considering the classical integral representation formula for the solutions to (1.1) (see e.g. [9, §II.2]).
In order to check (2), note that, since $B\neq 0$, the set $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$ contains at least a point different from 00. This implies that $\dim L\geq 1$. Denoting $N=\dim L$, consider $N$ points $q_{1},\ldots,q_{N}\in\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$ that are linearly independent (they exist because, otherwise, $L$ would not be the smallest linear space containing $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$) and let $C$ be the convex hull of the set $\{q_{1},-q_{1},q_{2},-q_{2},\ldots,q_{N},-q_{N}\}$. The interior of $C$ in the topology of $L$ is an open neighbourhood of the origin and it is contained in $\mathcal{C}_{\overline{q}=0}^{(A,B)p.c.}$ because of (1).
We now prove (3). First of all we claim that each vector field $X_{\alpha}$, defined in (4.1), is tangent to $L$ at all points of $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$. In order to check this, for what concerns a vector field $X_{\alpha_{o}}$, $\alpha_{o}\neq 0$, suppose that the claim is not true, i.e. that $X_{\alpha_{o}}|_{q}\notin L$ for some $q\in\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$. Since $X_{\alpha_{o}}$ has constant components, this implies that also $X_{\alpha_{o}}|_{\overline{q}=0}\notin L$. In this case, consider the differential problem $\dot{q}^{i}=-A^{i}_{j}q^{j}-\mu B_{\alpha_{o}}$, $q(0)=0$ for some small constant value $\mu>0$ and let $q(t)$ be a solution to this problem, hence a curve, which is tangent to the vector field $X_{0}(q)-\mu X_{\alpha_{o}}(q)$. Note that, in the linear system $(-A,B)$, this corresponds to the choice of the constant control $\mu e_{\alpha_{o}}$ which, for small $\mu>0$, belongs to $\mathscr{K}$ because it is a neighbourhood of the origin in $\mathbb{R}^{m}$, being $e_{\alpha_{o}}$ the suitable element of its canonical basis. Since $X_{\alpha_{o}}|_{\overline{q}=0}\notin L$ and $X_{0}|_{\overline{q}=0}=0$, for small values of $t$ the point $q(t)$ is not in $L$. This is however impossible because all points of such a solution $q(t)$ are in $\mathscr{O}^{(-A,-B)p.c.}_{\overline{q}=0}=\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}\subset L$. For what concerns the vector fields $X_{0}$ (which has not constant components), the proof is different. As before, assume that the claim does not hold for $X_{0}$, i.e. that there is a $\overline{q}^{\prime}\in\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$ such that $X_{0}|_{\overline{q}^{\prime}}$ is not tangent to $L$. We may therefore consider
  • (i)
    A solution $q(t)$ of a differential problem $\dot{q}=-Aq-Bu(t)$, $q(0)=0$, that ends at $\overline{q}^{\prime}$ and is determined by a piecewise constant control $u(t)$ (it exists because $\overline{q}^{\prime}\in\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}=\mathscr{O}^{(-A,-B)p.c.}_{\overline{q}=0}$);
  • (ii)
    A solution $q(t)$ to the differential problem $\dot{q}=-Aq$, $q(0)=\overline{q}^{\prime}$ (such a solution is an integral curve of the vector field $X_{0}$ and, since $X_{0}|_{\overline{q}^{\prime}}\notin L$, the points $q(t)$ are not in $L$ for small values of $t$; and again, it is an admissible curve for the controlled system $(-A,B)$ because $0\in\mathscr{K}$).
Composing the curves in (i) and (ii), we get a curve which is entirely included in $\mathscr{O}^{(-A,-B)p.c.}_{\overline{q}=0}=\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$, but also which is not entirely included in $L$ contradicting the assumption $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}\subset L$.
From the argument above, we see that all vector fields in the set $\mathscr{F}^{(A,B)}$ and, consequently, all vector fields in the spanned vector space $\widehat{\mathscr{F}^{(A,B)}}$ are tangent to $L$ at the points of an open (in the topology of $L$) neighbourhood $\mathscr{U}\subset L$ of 00, which is entirely contained in $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$ (it exists by (2)).Thus any integral curve $\sigma^{X}(s)$ of a vector field $X$ in $\widehat{\mathscr{F}^{(A,B)}}$ that starts from a point in $\mathscr{U}$ stays in $\mathscr{U}$ for small values of $s$. Consequently, for any $q\in\mathscr{U}\subset\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$, there exists an open neighbourhood of $q$ of the intrinsic topology of $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(q)$ which is included in $\mathscr{U}\subset L$. By Theorem 3.1, $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(\overline{q}_{o}=0)$ is an immersed submanifold of $\mathbb{R}^{n}$ of dimension greater than or equal to $\dim\operatorname{Lie}(\mathscr{F}^{(A,B)}_{\overline{q}_{o}=0})$. Thus:
  • $\dim\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(q_{o}=0)\leq\dim L$ because $\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(q_{o}=0)$ contains an open subset of its intrinsic topology included in $\mathscr{U}\subset L$;
  • $\dim\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(q_{o}=0)\geq\dim L$ because $\operatorname{Orb}^{\mathscr{F}}(q_{o}=0)$ contains $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$ which contains an open subset of $L$.
This implies that $\dim L=\dim\operatorname{Orb}^{\mathscr{F}^{(A,B)}}(\overline{q}_{o}=0)\geq\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{\overline{q}_{o}=0}$. ∎
We have now all ingredients to conclude the proof of the Kalman Theorem described at the beginning of this section. By (4.3) and the remarks just before that equality, we know that $\dim\operatorname{Kalman}^{(A,B)}=n$ if and only if $\dim\operatorname{Lie}(\mathscr{F}^{(A,B)})_{\overline{q}_{o}=0}=n$. Thus, by Proposition 5.1 (3), the linear subspace $L\subset\mathbb{R}^{n}$ of claim (2) of that proposition is equal to $\mathbb{R}^{n}$ and the set $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$ is a symmetric, convex neighbourhood of $\overline{q}=0$ in $\mathbb{R}^{n}$. In particular it is an open subset of $L=\mathbb{R}^{n}$. Thus if the Kalman condition holds, then $\mathcal{C}^{(A,B)p.c.}_{\overline{q}=0}$ contains a neighbourhood of $0_{\mathbb{R}^{n}}$ as we needed to prove.
Remark 5.2.
We point out that a generic problem of controllability is a “forward-in-time problem” in the following sense: It consists of the query of points that can be considered as starting points for the controlled evolutions that reach a desired destination in a finite time and moving just forward in time. On the other hand, the points of an orbit of $\mathscr{F}^{(A,B)}$, as considered in the Chow-Rashevskiĭ Theorem, are points, which can be reached moving both forward and backward in time along the evolutions. This is the crucial difference between the two problems. In the above proof, the argument where the“moving-just-forward-in-time” constraint on the evolutions (i.e. the main difference between the controllability problems and the Chow-Rashevskiĭ  type problems) is shown to be non effective for deriving the Kalman Theorem is when we prove that all the vector fields $X_{\alpha}$ (and in particular, the vector field $X_{0}$) are tangent to $L$. Indeed this is equivalent to the fact that $C_{\overline{q}=0}^{(A,B)p.c.}$ is a neighbourhood of 00 if and only if $C_{\overline{q}=0}^{(-A,B)p.c.}$ is so and hence if and only if $C_{\overline{q}=0}^{(\pm A,B)p.c.}$ is so as well. Note that $C_{\overline{q}=0}^{(\pm A,B)p.c.}$ is precisely the set of the controllable points of the problem given by the controlled dynamics $\dot{q}=\sigma Aq+Bu$ with controlling parameter $u=(u^{i})$ together with the additional discrete parameter $\sigma\in\{-1,1\}$.
6. Another proof of the Kalman Theorem, based on general facts on flows of vector fields
We now give our second proof of the Kalman Theorem or, more precisely, only of the sufficiency part of that theorem. In fact we already pointed out that the necessity part is just a consequence of a linear algebra argument, namely the same that proves the necessity part in the corollary of the Chow-Rashevskiĭ Theorem that we are considering in this paper.
This second proof is based on a translation of the control problem given by a linear system $(A,B)$ into a corresponding equivalent problem on certain distinguished curves and flows of certain vector fields in the extended space-time =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$. We introduce these special curves and vector fields and the above described translation in the next subsections §6.1 and §6.2. The actual proof is in the concluding subsection §6.3.
6.1. Stepped graphs in the extended space-time
Assume that $\mathscr{K}\subset\mathbb{R}^{m}$ is an open neighborhood of the origin. Consider now the manifold =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}=\mathbb{R}^{1+n}\times\mathscr{K}$ (which is an open subset of $\mathbb{R}^{1+n+m}$) and denote by $t$, $q=(q^{i})$ and $u=(u^{a})$ the standard cartesian coordinates of $\mathbb{R}$, $\mathscr{Q}$ and $\mathscr{K}\subset\mathbb{R}^{m}$, respectively, so that each $x\in\mathscr{M}$ is determined by a tuple $(t,q^{i},u^{a})$. As we mentioned in the Introduction, we name $\mathscr{M}$ the extended space-time associated with the system (1.1) and we denote by
$\pi^{\mathbb{R}}:\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}\to\mathbb{R}\ ,\ \ \pi^{\mathscr{Q}}:\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}\to\mathscr{Q}\ ,\ \ \pi^{\mathscr{K}}:\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}\to\mathscr{K}$
the three standard projections of the cartesian product $\mathscr{M}$ onto its three factors.
A stepped (completed) graph of a solution of (1.1) is a piecewise regular oriented curve (see §2.3) $\eta=\eta_{1}\ast\eta_{2}\ast\ldots\ast\eta_{2r+1}$ in $\mathscr{M}$, obtained by composing an odd number of regular arcs and satisfying the following three conditions (Fig.1):
  • (1)
    Each odd arc $\eta_{2k+1}$, $0\leq k\leq r$, is the graph
    $\eta_{2k+1}=\{\ (t,q_{k}(t),u_{k}(t))\ ,\ t\in[t_{k},t_{k}+\sigma_{k}]\ \}$
    of a map $t\longmapsto(q_{k}(t),u_{k}(t))\in\mathscr{Q}\times\mathscr{K}$, $t\in[t_{k},t_{k}+\sigma_{k}]$, in which $u_{k}(t)\equiv c_{k}$ is constant and $q_{k}(t)$ is the corresponding solution to (1.1); we denote its initial point by $\widetilde{q}_{k}=q_{k}(t_{k})$;
  • (2)
    For any odd regular arc $\eta_{2k+1}$ with $k\geq 1$, the initial condition $\widetilde{q}_{k}$ of $q_{k}(t)$ is equal to the final point $\widehat{q}_{k-1}:=q_{k-1}(t_{k-1}+\sigma_{k-1})$ of the solution $q_{k-1}(t)$;
  • (3)
    Each even regular arc $\eta_{2(k+1)}$, $10\leq k\leq r-$ 1, is the segment in ×$\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$ that joins the final point of $\eta_{2k+1}$ to the initial point of $\eta_{2(k+1)+1}$.
$t$$w$$q$$(t_{0},q_{0},w_{0})$
Fig. 1 – A stepped graph in =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$
Note that any stepped graph $\eta=\eta_{1}\ast\eta_{2}\ast\ldots\ast\eta_{2r+1}$ of a solution satisfies the following:
  • The projection on ×$\mathbb{R}\times\mathscr{Q}$ of each even arc $\eta_{2(k+1)}$ is a singleton, i.e. the set which contains just the point given by the projection onto ×$\mathbb{R}\times\mathscr{Q}$ of the final point of $\eta_{2k+1}$ (which is also the initial point of $\eta_{2k+3})$
    $\big{\{}(t_{k-1}+\sigma_{k-1},\widehat{q}_{k-1})\big{\}}=\big{\{}(t_{k},\widetilde{q}_{k})\big{\}}\ ;$
  • The projection of $\eta$ onto ×$\mathbb{R}\times\mathscr{Q}$ coincides with the graph of the solution $t\to q(t)$ to (1.1) with initial condition $\widetilde{q}_{0}$ and the piecewise constant control curve $u(t)$, which is equal to the constants $c_{0},c_{1},\ldots,c_{r}$ along the time intervals $(t_{k},t_{k}+\sigma_{k})\subset[0,T]$ (Fig.2);
    $t$$q$$(t_{0},q_{0})$
    Fig. 2 – The graph of a solution in ×$\mathbb{R}\times\mathscr{Q}$ corresponding to a stepped graph
  • $\eta$ is determined by its odd regular arcs (the even arcs are just segments and are completely determined by the endpoints of the odd arcs).
Conversely, if $u:[0,T]\to K\subset\mathbb{R}^{m}$ is a piecewise constant control curve in $\mathscr{K}\subset\mathbb{R}^{m}$ and takes constant values $c_{0},c_{1},\ldots,c_{r}\in\mathscr{K}$ on the subintervals
$[0,t_{1})\ ,\ (t_{1},t_{2})\ ,\ldots\ ,\ (t_{r},T]\qquad\text{with}\qquad[0,T]=\bigcup_{j=0}^{r}[t_{j},t_{j+1}]$
and denoting by $q(t)$, $t\in[0,T]$, the solution to (1.1) corresponding to $u(t)$ and the initial value $q(0)=\widetilde{q}_{0}$, there is just one stepped graph $\eta=\eta_{1}\ast\eta_{2}\ast\ldots\ast\eta_{2r+1}$ in $\mathscr{M}$, whose odd regular arcs $\eta_{2k+1}$ are determined by the restrictions $q|_{[(t_{k},t_{k+1}]}$, $0\leq k\leq r$. The projection on ×$\mathbb{R}\times\mathscr{Q}$ of $\eta$ is the graph $\{(t,q(t))\ ,\ t\in[0,T]\}\subset\mathbb{R}\times\mathscr{Q}$ of such a solution.
Given a point $\overline{x}=(\overline{t},\overline{q},\overline{w})\in\mathscr{M}$, the $\mathscr{M}$-attainable set of $\overline{x}$ in the time $T$ is the set
-$\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}}{=}\bigg{\{}y\in\mathscr{M}\,:\,y\ \text{is the final point of a stepped graph}\ \eta\\ \text{that starts from $\overline{x}$ and is such that}\ \pi^{\mathbb{R}}(\eta)=[0,T]\bigg{\}}\ .$ (6.1)
From the previous discussion and (3.5), it follows that
$\mathcal{C}_{\overline{q}}^{(A,B)p.c.}=\mathscr{O}_{\overline{q}}^{(\mathbb{A},\mathbb{B})p.c.}=\bigcup_{T>0,\overline{w}\in\mathscr{K}}\pi^{\mathscr{Q}}\left(\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}:=(0,\overline{q},\overline{w})}\right)\ .$ (6.2)
Remark 6.1.
Due to the above relation, whenever there exists a point $\overline{w}\in\mathscr{K}$ and a time $T>0$, such that -$\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}:=(0,0,\overline{w})}$ projects onto a neighbourhood of $0_{\mathbb{R}^{n}}$ via $\pi^{\mathscr{Q}}$, it follows immediately that $\mathcal{C}_{\overline{q}}^{(A,B)p.c.}=\mathscr{O}_{\overline{q}}^{(\mathbb{A},\mathbb{B})p.c.}$ contains a neighbourhood of the origin.
Our second proof of the sufficiency part of the Kalman Theorem consists precisely in exploiting Remark 6.1, namely in showing that the Kalman condition implies that, for $\overline{w}=0$, the projection of -$\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}:=(0,0,0)}$ onto $\mathscr{Q}$ has the desired property for any choice of $T>0$.
6.2. Final points of stepped graphs as points in orbits of flows of vector fields
Consider a stepped graph $\eta=\eta_{1}\ast\eta_{2}\ast\ldots\ast\eta_{2r+1}$ in =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$, corresponding to a solution to (1.1) for a piecewise constant control $u:[0,T]\to\mathscr{K}$ with values $c_{k}\in\mathscr{K}$, $0\leq k\leq r$, on time intervals $[t_{k},t_{k}+\sigma_{k}]$, with $t_{0}=0$ and $t_{k}:=t_{k-1}+\sigma_{k-1}$. Let also $\overline{x}$ and $\overline{y}$ be the initial and final points of $\eta$.
We observe that each odd arc $\eta_{2k+1}$, $0\leq k\leq r$, is an integral curve of the vector field
$\mathbb{T}=\frac{\partial}{\partial t}+\left(\mathbb{A}^{i}_{j}q^{j}+\mathbb{B}^{i}_{a}u^{a}\right)\frac{\partial}{\partial q^{i}}=\frac{\partial}{\partial t}-\left(A^{i}_{j}q^{j}+B^{i}_{a}u^{a}\right)\frac{\partial}{\partial q^{i}}\ ,$
while each even arc $\eta_{2(k+1)}$ (that is, each of the segments that join the points $(t_{k+1},\widehat{q}_{k},c_{k})$ to $(t_{k+1},\widetilde{q}_{k+1}=\widehat{q}_{k},c_{k+1})$) are integral curves, with integration parameter running in $[0,1]$, of the vector fields
$\lambda_{(k)}^{a}\frac{\partial}{\partial u^{a}}\ ,\qquad\lambda_{(k)}^{a}:=c_{k}^{a}-c_{k-1}^{a}\ ,\qquad 1\leq k\leq r\ .$
Setting $W_{a}:=\frac{\partial}{\partial u^{a}}$, we conclude that the final point $\overline{y}$ is equal to
$\overline{y}=\Phi^{\mathbb{T}}_{\sigma_{r}}\circ\Phi_{1}^{\lambda_{(r)}^{a}W_{a}}\circ\Phi^{\mathbb{T}}_{\sigma_{r}}\circ\ldots\circ\Phi_{1}^{\lambda_{(2)}^{a}W_{a}}\circ\Phi^{\mathbb{T}}_{\sigma_{1}}\circ\Phi_{1}^{\lambda_{(1)}^{a}W_{a}}\circ\Phi^{\mathbb{T}}_{\sigma_{0}}(\overline{x})\ .$ (6.3)
The next lemma shows that such expression is equivalent to another, much more convenient for our purposes. In its statement, for any $s\in\mathbb{R}$ and any vector field $X$ on =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$, we adopt the notation $X^{[s]}$ to indicate the vector field defined by
$X^{[s]}=\Phi^{\mathbb{T}}_{s*}(X)\ ,$
that is the push-forward of $X$ by the diffeomorphism $\Phi^{\mathbb{T}}_{s}$. We recall that such a pushed-forward vector field $X^{[s]}$ is the unique vector field whose integral curves are the curves in $\mathscr{M}$ defined by
[…]
Lemma 6.2.
Setting $\tau_{k}:=\sum_{j=k}^{r}\sigma_{j}$, $1\leq k\leq r$, the point $\overline{y}$ in (6.3) is also equal to the point
$\overline{y}=\big{(}\Phi_{1}^{\lambda_{(r)}^{a}W^{[\tau_{r}]}_{a}}\circ\Phi_{1}^{\lambda_{(r-1)}^{a}W^{[\tau_{r-1}]}_{a}}\circ\ldots\ \circ\Phi_{1}^{\lambda_{(1)}^{a}W^{[\tau_{1}]}_{a}}\big{)}(\overline{x}^{\prime})\ \ \text{where}\ \overline{x}^{\prime}:=\Phi^{\mathbb{T}}_{T}(\overline{x})\ .$ (6.4)
In particular, if $\overline{x}=(0,\overline{q}=0,\overline{w}=0)$,
$\overline{y}=\big{(}\Phi_{1}^{\lambda_{(r)}^{a}W^{[\tau_{r}]}_{a}}\circ\Phi_{1}^{\lambda_{(r-1)}^{a}W^{[\tau_{r-1}]}_{a}}\circ\ldots\ \circ\Phi_{1}^{\lambda_{(1)}^{a}W^{[\tau_{1}]}_{a}}\big{)}(\overline{x})\ .$ (6.5)
$t$$w$$q$Curve $s\mapsto\Phi^{\mathbb{T}}_{s}(\overline{x})$, $s\in[0,\sigma_{1}+\ldots+\sigma_{r}]=[0,T]$$\overline{x}^{\prime}$$\overline{y}$ $\overline{x}$urves $s\mapsto\Phi_{s}^{\lambda_{(j)}^{a}W_{a}}$urves $s\mapsto\Phi_{s}^{\lambda_{(j)}^{a}W^{[\tau_{j}]}_{a}}$
Fig. 3 – The black curve from $\overline{x}^{\prime}$ to $\overline{y}$ is obtained changing the
composition of flows (6.3) into a new one by an iterated application of (6.6).
Proof.  By the identity (2.2), any composition of flows of the form $\Phi^{\mathbb{T}}_{s}\circ\Phi_{t}^{X}$ can be replaced by
[…]
Applying this identity iteratively, after a finite number of steps formula (6.3) takes the form (see Fig.3)
$\overline{y}=\left(\Phi_{1}^{\lambda_{(r)}^{a}W^{[\sigma_{r}]}_{a}}\circ\Phi_{1}^{\lambda_{(r-1)}^{a}W^{[\sigma_{r-1}+\sigma_{r}]}_{a}}\circ\ldots\ \circ\Phi_{1}^{\lambda_{(1)}^{a}W^{[\sigma_{r}+\sigma_{r-1}+\ldots\sigma_{1}]}_{a}}\right)\circ\\ \circ\Phi^{\mathbb{T}}_{\sigma_{r}+\ldots+\sigma_{1}+\sigma_{0}}(\overline{x})\ .$ (6.7)
Since $\sigma_{r}+\ldots+\sigma_{1}+\sigma_{0}=T$, (6.4) follows. The last claim is an immediate consequence of (6.4) and the fact that $\Phi^{\mathbb{T}}_{t}(0)=0$ for all $t\in\mathbb{R}$.
From Lemma 6.2 if follows that if $\overline{x}_{o}=(0,0,0)$, a point $\overline{y}$ belongs to -$\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}_{o}}$ for some $T>0$ if and only if it has the form (6.5) for a finite set of constants $\lambda_{(k)}=c_{k}-c_{k-1}$, $1\leq k\leq r$ (which correspond to constant controls $c_{k}\in\mathscr{K}\subset\mathbb{R}^{m}$) and a finite sequence of real numbers $\tau_{k}$ satisfying the inequalities $T>\tau_{1}>\tau_{2}>\ldots>\tau_{r}>0$.
6.3. Our second proof of the sufficiency part of the Kalman Theorem
Consider the constant vector fields on $\mathscr{M}$ defined recursively by
$\begin{split}&W^{(0)}_{1}:=W_{1}=\frac{\partial}{\partial u^{1}}\ ,\quad W^{(0)}_{2}:=W_{2}=\frac{\partial}{\partial u^{2}}\ ,\qquad\ldots\ ,\qquad W_{m}^{(0)}:=W_{m}=\frac{\partial}{\partial u^{m}}\ ,\\ &W_{1}^{(1)}:=[\mathbb{T},W_{1}]=B_{1}^{i}\frac{\partial}{\partial q^{i}}\ ,\quad W_{2}^{(1)}:=[\mathbb{T},W_{2}]=B_{2}^{i}\frac{\partial}{\partial q^{i}}\ ,\ \ldots\ \\ &\hskip 199.16928pt\ldots\ \ W^{(1)}_{m}:=[\mathbb{T},W_{m}]=B_{m}^{i}\frac{\partial}{\partial q^{i}}\ ,\\ &W_{1}^{(2)}:=[\mathbb{T},W^{(1)}_{1}]=A_{j}^{i}B_{1}^{j}\frac{\partial}{\partial q^{i}}\ ,\quad W_{2}^{(2)}:=[\mathbb{T},W^{(1)}_{2}]=A_{j}^{i}B_{2}^{j}\frac{\partial}{\partial q^{i}}\ ,\ \ldots\\ &\hskip 199.16928pt\ldots\ ,\ W^{(2)}_{m}:=[\mathbb{T},W^{(1)}_{m}]=A^{i}_{j}B_{m}^{j}\frac{\partial}{\partial q^{i}}\ ,\\ &\hskip 170.71652pt\vdots\\ &W_{1}^{(\ell)}:=[\mathbb{T},W^{(\ell-1)}_{1}]=A_{j_{1}}^{i}A_{j_{2}}^{j_{1}}\ldots A^{j_{\ell-2}}_{j_{\ell-1}}B_{1}^{j_{\ell-1}}\frac{\partial}{\partial q^{i}}\ ,\ \ldots\\ &\hskip 113.81102pt\ldots\ ,\ W^{(\ell)}_{m}:=[\mathbb{T},W^{(\ell-1)}_{m}]=A_{j_{1}}^{i}A_{j_{2}}^{j_{1}}\ldots A^{j_{\ell-2}}_{j_{\ell-1}}B_{m}^{j_{\ell-1}}\frac{\partial}{\partial q^{i}}\ ,\\ &\hskip 170.71652pt\vdots\end{split}$ (6.8)
Note that the $i-$th component of each vector field $W^{(\ell)}_{a}$, $1\leq a\leq m$, is always equal to the $i-$th component of the vector $\underset{(\ell-1)-\text{times}}{\underbrace{A{\cdot}\ldots{\cdot}A}}{\cdot}B_{a}$.
Moreover, for any integer $\ell\geq 0$, let us denote by $n_{\ell}$ the dimension of the vector space spanned by the $m{\cdot}(\ell+1)$ vector fields $W_{a}^{(k)}$, $1\leq a\leq m$, $0\leq k\leq\ell$ at a point $x_{o}\in\mathscr{M}$:
[…]
(since all considered vector fields are constant, $n_{\ell}$ is independent of $x_{o}$). Clearly the sequence of integers $n_{0},n_{1},n_{2},\ldots$ is non-decreasing and each of them is less than or equal to $m+n=\dim\mathscr{K}+\dim\mathscr{Q}$. We denote by $\ell_{\text{max}}$ the smallest integer for which $n_{\ell_{\text{max}}}$ is maximal. One can immediately realise that (see (1.2))
$n_{\ell_{\max}}=m+\dim\operatorname{Kalman}^{(A,B)}$ (6.9)
and that $n_{\ell_{\max}}=m+n$ if and only if $\dim\operatorname{Kalman}^{(A,B)}=n$.
Let us now fix an ordered tuple $\mathcal{B}=(E_{1},\ldots,E_{n_{\ell_{\max}}})$, in which each vector field $E_{A}$ is a vector belonging to the set $\{W_{a}^{(\ell)}\}_{\begin{smallmatrix}1\leq a\leq m\\ 0\leq\ell\leq\ell_{\max}\end{smallmatrix}}$, and such that they constitute a basis for the following vector space at $x_{o}\in\mathscr{M}$ (and thus at any other point):
[…]
Since $n_{\ell_{{\max}}}$ is the maximal dimension, each vector field $W^{(\ell)}_{a}$, $\ell\geq\ell_{\max}+1$, admits a unique expansion with constant coefficients in terms of the vector fields $E_{A}$, $1\leq A\leq n_{\ell_{{\max}}}$. Since each $E_{A}$ has the form $W_{b}^{(\ell)}$ for some $1\leq b\leq m$ and $0\leq\ell\leq\ell_{\max}$, we have that all vector fields $W^{(r)}_{a}$, $r\geq\ell_{\max}+1$, admits a unique expansion of the form
$W^{(r)}_{a}=\sum_{\begin{smallmatrix}0\leq\ell\leq\ell_{\max}\\ 1\leq b\leq m\end{smallmatrix}}\lambda^{r|b}_{a|\ell}W^{(\ell)}_{b}\ ,\qquad r\geq\ell_{{\max}}+1\ ,$ (6.11)
where each coefficient $\lambda^{r|b}_{a|\ell}$ is either 00 or equal to the component of $W^{(r)}_{a}$ in the direction of $W^{(\ell)}_{b}$ in case $W^{(\ell)}_{b}$ is one of the chosen basis $\mathcal{B}=(E_{A})$.
Lemma 6.3.
For any $1\leq a\leq m$ and $s\in\mathbb{R}$, the vector field $W_{a}^{[s]}=\Phi^{\mathbb{T}}_{s*}(W_{a})$ has a unique expansion of the form
$W_{a}^{[s]}=\sum_{\ell=0}^{\ell_{{\max}}}\left((-1)^{\ell}\frac{s^{\ell}}{\ell!}\delta_{a}^{b}+s^{\ell_{{\max}}+1}D_{a|\ell}^{b}(s)\right)W_{b}^{(\ell)}\ ,$ (6.12)
where $D_{a|\ell}^{b}(s)$ is the sum of the converging series
$D_{a|\ell}^{b}(s):=(-1)^{\ell_{\max}+1}\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{s^{\ell}}{(\ell+\ell_{{\max}}+1)!}\lambda_{a|\ell}^{\ell+\ell_{{\max}}+1|b}$ (6.13)
and $\delta_{a}^{b}$ is the Kronecker delta. In particular,
  • for each $s\in\mathbb{R}$, the vector field $W_{a}^{[s]}$ is a finite linear combination of the vector fields $W_{b}^{(\ell)}$, $1\leq b\leq m$, $0\leq\ell\leq\ell_{\max}$, and
  • for each $s,s^{\prime}\in\mathbb{R}$, the diffeomorphism $\Phi^{W^{[s]}_{a}}_{s^{\prime}}$ maps the hypersurfaces $\{t=\text{const.}\}$. } into themselves.
Proof.  We recall that for any $s\in\mathbb{R}$ and any vector field $X$ on $\mathscr{M}$,
$\frac{d}{ds}\Phi^{\mathbb{T}}_{s_{*}}(X)\bigg{|}_{s}=\frac{d}{dh}\Phi^{\mathbb{T}}_{h_{*}}(X^{[s]})\bigg{|}_{h=0}=\frac{d}{dh}\Phi^{-\mathbb{T}}_{h}{}^{*}(X^{[s]})\bigg{|}_{h=0}=-[\mathbb{T},X^{[s]}]\ .$ (6.14)
Due to this, for any $a\in\{1,\ldots,m\}$, the one-parameter family of vector fields $s\mapsto W_{a}^{[s]}$, $s\in\mathbb{R}$, is the unique solution to the differential problem
$\frac{dW_{a}^{[s]}}{ds}\bigg{|}_{s}=-[\mathbb{T},W_{a}^{[s]}]\ ,\qquad W_{a}^{[s=0]}=W_{a}\ .$ (6.15)
On the other hand, if $W_{a}^{(\ell)i}$, $W_{a}^{(\ell)b}$ denote the components of the vector field $W_{a}^{(\ell)}=W_{a}^{(\ell)i}\frac{\partial}{\partial q^{i}}+W_{a}^{(\ell)b}\frac{\partial}{\partial u^{b}}$, using the iterative definition (6.8) of such vector field, one can directly check that the power series
$X^{i}(s)=\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{s^{\ell}}{\ell!}W_{a}^{(\ell)i}\ ,\qquad X^{b}(s):=\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{s^{\ell}}{\ell!}W_{a}^{(\ell)b}$
uniformly converges on any closed interval $[s_{1},s_{2}]\subset\mathbb{R}$. This implies that the vector field
$X(s)=X^{i}(s)\frac{\partial}{\partial q^{i}}+X^{b}(s)\frac{\partial}{\partial u^{b}}$
is well defined and real analytic for any $s\in\mathbb{R}$, and it is such that $X(0)=W_{a}$ and $\frac{dX(s)}{ds}\big{|}_{s}=-[\mathbb{T},X(s)]$. From uniqueness of the solution to (6.15), we conclude that $W^{[s]}_{a}=X(s)=\sum_{\ell=0}^{\infty}(-1)^{\ell}\frac{s^{\ell}}{\ell!}W_{a}^{(\ell)}$. We now recall that each vector field $W_{a}^{(\ell)}$, $\ell\geq\ell_{{\max}}+1$, $1\leq a\leq m$, admits a unique expansion (6.11). Replacing such expansion in all terms corresponding to the powers $s^{\ell}$, $\ell\geq\ell_{{\max}}+1$, we get (6.12). The last two claims are consequences of (6.12) and the fact that each vector field $W_{b}^{(\ell)}$ has trivial component along $\frac{\partial}{\partial t}$.
For any $\varepsilon>0$ and any ordered pair of integers $(a,\ell)$, $1\leq a\leq m$, $0\leq\ell\leq\ell_{{\max}}$, let
$\tau_{a,\ell}:=\frac{a+\ell m}{m\cdot(\ell_{{\max}}+1)}\ ,\qquad\mathbb{W}_{a}^{(\ell)}:=\ell!W_{a}^{[\varepsilon\tau_{a,\ell}]}\ .$
Note that each real number $\varepsilon\tau_{a,\ell}$ belongs to the interval $\left[\frac{1}{m(\ell_{{\max}}+1)}\varepsilon,\varepsilon\right]$ and the latter is a subinterval of $(0,T)$ for a prescribed value of $T$, provided that $\varepsilon$ is sufficiently small. Imposing the lexicographic order for the set of pairs $(a,\ell)$ and denoting each such pair by a single index $1\leq A\leq m\cdot(\ell_{{\max}}+1)$, we may consider the tuples of vector fields
$\{\mathbb{W}_{A}\}_{1\leq A\leq m\ell_{{\max}}}=\{\mathbb{W}_{a}^{(\ell)}\}_{\begin{smallmatrix}1\leq a\leq m\\ 0\leq\ell\leq\ell_{{\max}}\end{smallmatrix}}\ ,\qquad\{W_{B}\}_{1\leq B\leq m\ell_{{\max}}}=\{W_{a}^{(\ell)}\}_{1\leq a\leq m,0\leq\ell\leq\ell_{{\max}}}$
and the squared matrix $\mathscr{A}=\big{(}\mathscr{A}_{A}^{B}\big{)}$, with entries given by the components of the vector fields $\mathbb{W}_{A}$ in terms of the vector fields $W_{B}$ according to (6.12). Notice that $\mathscr{A}$ has the form $\mathscr{A}=\widehat{\mathscr{A}}+O(\varepsilon^{\ell_{{\max}}+1})$, where
[…]
and $O(\varepsilon^{\ell_{{\max}}+1})$ stands for a squared matrix whose entries are infinitesimal of order $\ell_{{\max}}+1$ with respect to $\varepsilon$.
Lemma 6.4.
For any $T>0$, there exists $\varepsilon\in\left(0,\frac{T}{2}\right)$ such that the matrix $\mathscr{A}$ is invertible.
Proof.  Consider the matrix
$\mathcal{B}=\left(\begin{array}[]{cccc}\mathcal{B}_{1}&0&\dots&0\\ 0&\mathcal{B}_{2}&\dots&0\\ \vdots&\vdots&\ddots&0\\ 0&0&\dots&\mathcal{B}_{m}\end{array}\right)\ ,\ \ \mathcal{B}_{a}:=\left(\begin{array}[]{ccccc }1&\tau_{a,0}&\tau^{2}_{a,0}&\ldots&\tau_{a,0}^{\ell_{{\max}}}\\ 1&\tau_{a,1}&\tau^{2}_{a,1}&\ldots&\tau_{a,1}^{\ell_{{\max}}}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&\tau_{a,\ell_{{\max}}}&\tau^{2}_{a,\ell_{{\max}}}&\ldots&\tau_{a,\ell_{{\max}}}^{\ell_{{\max}}}\end{array}\right)\ .$ (6.17)
Factoring out an appropriate power $\varepsilon^{r}$, $0\leq r\leq\ell_{{\max}}$, from the terms of each column of $\mathscr{A}$, we get that
$\det\mathscr{A}=\left(\prod_{0\leq r\leq\ell_{\max}}\varepsilon^{r}\right)^{m}\det\left(\mathcal{B}+O(\varepsilon)\right)\ ,$
where $O(\varepsilon)$ denotes a matrix whose entries are infinitesimal of the same order of $\varepsilon$. Therefore, by continuity and the well-known formula for the determinants of Vandermonde matrices (see e.g. [7]),
$\det\mathscr{A}(\varepsilon)=\left(\prod_{0\leq r\leq\ell_{\max}}\varepsilon^{r}\right)^{m}\prod_{1\leq a\leq m}\left(\prod_{1\leq\ell<\ell^{\prime}\leq\ell_{{\max}}}(\tau_{a,\ell^{\prime}}-\tau_{a,\ell})+O(\varepsilon)\right)$
where now $O(\varepsilon)$ stands for a function which is an infinitesimal of the same order of $\varepsilon$. Since each factor $\prod_{1\leq\ell<\ell^{\prime}\leq\ell_{{\max}}}(\tau_{a,\ell^{\prime}}-\tau_{a,\ell})+O(\varepsilon)$ tends to a non-zero value for $\varepsilon$ tending to 00, the lemma follows.
Theorem 6.5.
Assume that $\mathscr{K}\subset\mathbb{R}^{m}$ is an open set and that $\mathcal{C}_{\overline{q}}^{(A,B)p.c.}$ contains a point $q$ that reaches $\overline{q}$ in a time $T>0$ through a solution $q^{(q,u_{o})}$ with $u_{o}(t)$ piecewise constant. If $\operatorname{Kalman}^{(A,B)}=n$, then $\mathcal{C}_{\overline{q}}^{(A,B)p.c.}$ contains a neighbourhood of the point $q$.
[…]
Let $q$, $\mathcal{C}_{\overline{q}}^{(A,B)p.c.}$ and $u_{o}(t)$ as in the hypothesis and assume that $u_{o}(t)$ is constant. Pick $\varepsilon\in(0,T)$ and consider the map
$f^{(\varepsilon)}:\mathbb{R}^{m{\cdot}(\ell_{{\max}}+1)}\longrightarrow\mathscr{M}\\ f^{(\varepsilon)}(s_{1,0},s_{2,0},\ldots,s_{a,\ell},\ldots,s_{m,\ell_{{\max}}}):=\bigg{(}\Phi_{s_{1,0}}^{W_{1}^{[\varepsilon\tau_{1,0}])}}\circ\Phi_{s_{2,0}}^{W_{2}^{[\varepsilon\tau_{2,0}]}}\circ\ldots\circ\Phi_{s_{a,\ell}}^{W_{a}^{[\varepsilon\tau_{a,\ell}]}}\circ\ldots\\ \ldots\circ\Phi_{s_{m,\ell_{{\max}}}}^{W_{m-1}^{[\varepsilon\tau_{m-1,\ell_{{\max}}}]}}\circ\Phi_{s_{m,\ell_{{\max}}}}^{W_{m}^{[\varepsilon\tau_{m,\ell_{{\max}}}]}}\circ\Phi_{T}^{\mathbb{T}}\bigg{)}(0,\overline{q},c)\ .$ (6.18)
By construction and Lemma 6.3, the following properties hold:
  • (i)
    $f^{(\varepsilon)}(0)=\Phi_{T}^{\mathbb{T}}(0,\overline{q},c)=(T,q,c)$ and for any $s\in\mathbb{R}^{m{\cdot}(\ell_{{\max}}+1)}$, the point $f^{(\varepsilon)}(s)$ is in the hypersurface $\{t=T\}\subset\mathscr{M}$;
  • (ii)
    By Lemma 6.2, each point of the form $\overline{y}=f^{(\varepsilon)}(s)$ belongs to -$\mathscr{M}\text{\it-Att}^{(T)}_{\overline{x}_{o}}$, $x_{o}=(0,\overline{q},c)$;
  • (iii)
    The columns of the Jacobian $Jf^{(\varepsilon)}\big{|}_{0}$ are the components of the vector fields
    $\frac{df^{(\varepsilon)}(0,0,\ldots,0,s_{a,\ell}=t,0,\ldots,0)}{dt}\bigg{|}_{t=0}=\ell!W_{a}^{[\varepsilon\tau_{a,\ell}]}=\mathbb{W}_{a}^{(\ell)}\ \ .$ (6.19)
By iii) and Lemma 6.4, for any sufficiently small $\varepsilon>0$, the rank of $Jf^{(\varepsilon)}\big{|}_{0}$ is equal to the dimension of the vector space spanned by the vector fields $W_{a}^{(\ell)}$, i.e. $m+\dim\operatorname{Kalman}^{(A,B)}$. Such a rank is equal to $m+n=\dim\mathscr{K}+\dim\mathscr{Q}$ if and only if $\dim\operatorname{Kalman}^{(A,B)}=n$. If this is the case, by the Inverse Function Theorem the image of $f^{(\varepsilon)}$ contains a neighbourhood of the point $y_{o}=(T,q,c)$ in the hypersurface $\{t=T\}\subset\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$. Since the map $\pi^{\mathscr{Q}}:\mathscr{M}\to\mathscr{Q}$ has maximal rank at all points of $\{t=T\}$, such a neighbourhood of $y_{o}$ projects onto an open neighbourhood of $q$ and, by (6.2), such an open set is a neighbourhood of $q$ which is entirely included in $\mathscr{O}_{\overline{q}}^{(\mathbb{A},\mathbb{B})p.c.}=\mathcal{C}_{\overline{q}}^{(A,B)p.c.}$.
Our second proof of the sufficiency part of the Kalman Theorem, described at the beginning of this section, is now a trivial consequence of Theorem 6.5: It suffices to observe that such a theorem applies to the case in which $q=0_{\mathbb{R}^{n}}$ and $\bar{q}=0_{\mathbb{R}^{n}}$, because the origin can be considered as joined to itself by means the trivial solution of (1.1), determined by the constant control $u_{o}(t)\equiv 0$, which is an admissible control by the hypothesis of Theorem 3.2.
Declarations
Ethical Approval: Not applicable.
Funding: Not apllicable.
Acknowledgments: F. Bagagiolo is a member of INDAM-Gruppo Nazionale per l’Analisi Matematica la Probabilità e le loro Applicazioni; M. Zoppello is a member of INDAM-Gruppo Nazionale di Fisica Matematica. F. Bagagiolo thanks Paolo Venturini for discussions on some parts of the subjects.